OpenStudy (loser66):

Use a double integral in polar coordinates to find the volume of the solid bounded by the elliptical paraboloids z =x^2+y^2 and z=6-(x^2+2y^2) please, help

OpenStudy (loser66):

@reemii

OpenStudy (loser66):

My question: how to find the polar coordinate of an ellipse? We know that if it is a circle, then polar coordinates are x = r cos t, y = r sint, how about ellipse?

OpenStudy (irishboy123):

in cylindrical \(z = r^2\) \(z = 6 - r^2\)

OpenStudy (irishboy123):

\(V = \int_V r dr d\theta dz\)

OpenStudy (loser66):

It is not cylindrical, the intersection of the surface is an ellipse, right?

OpenStudy (irishboy123):